Let 'a' and 'b' be relatively prime integers. Find all values of (a + 2b, a − 2b).
1
Expert's answer
2019-10-18T09:24:27-0400

Given that "a" and "b" are relatively prime integers. Then "gcd(a,\ b)=1" or in equivalent notation "(a,\ b)=1."

Let "d=(a+2b, \ a-2b)." Then it follows that "d|(a+2b)" and "d|(a-2b)."

That is "d|(m(a+2b)+n(a-2b)),"

"2(a+2b)+2(a-2b)=4a" and thus "d|4a,"

"2(a+2b)-2(a-2b)=4b" and thus "d|4b."

Hence "d|(4a,\ 4b)". We have that "d|4."

We can rule out every possibility except "d=1,2" or "d=4." However, we do not yet know which of those three values are actually possible.

Let's show that all possible values of "d\in \lbrace 1,2,4\rbrace" actually occur:

"a=1,b=0" gives "d=1,"

"a=0, b=1" gives "d=2,"

"a=2, b=1" gives "d=4."

Therefore, the answer is "d\in \lbrace 1,2,4\rbrace."